satellite73 (satellite73):

prove that given f, g differentiable blah blah that \[(fg)'=f'g+g'f\] using only that \[(f^2)'=2ff'\]

myininaya (myininaya):

so why can't we use definition of a derivative lol

OpenStudy (anonymous):

you can but then you have to do that stupid non intuitive add and subtract business, when in fact it is only algebra

myininaya (myininaya):

because i'm only allowed to use (f^2)'=2ff'

OpenStudy (anonymous):

yes only the rule for squares which easy to prove.

myininaya (myininaya):

so we have (f^2)'=2ff'=ff'+ff'... so im thinking

OpenStudy (zarkon):

maybe look at ((f+g)^2)'

myininaya (myininaya):

can we use (f+g)'=f'+g'?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok well then its done

OpenStudy (zarkon):

nice way of doing it

OpenStudy (anonymous):

i thought so because the usual method makes more out of it than there really is

myininaya (myininaya):

[(f+g)^2]' =2(f+g)(f+g)' =2(f+g)(f'+g') =2(ff'+fg'+gf'+gg') =2ff'+2fg'+2gf'+2gg' =(f^2)'+2fg'+2gf'+(g^2)' [(f+g)^2]'-(f^2)'-(g^2)'=2(fg'+gf') [(f+g)^2-f^2-g^2]'=2(fg'+gf') [f^2+2fg+g^2-f^2-g^2]'=2(fg'+gf') [2fg]'=2(fg'+gf') (fg)'=fg'+gf'

myininaya (myininaya):

i like this satellite

OpenStudy (anonymous):

lol i like it too

OpenStudy (anonymous):

usually start with \[(f+g)^2=f^2+g^2+2fg\] \[fg=\frac{1}{2}((f+g)^2-f^2-g^2)\] which also says that any product can be written as the sum and difference of squares

myininaya (myininaya):

i never thought to prove the product rule this way

myininaya (myininaya):

thanks for new knowledge

OpenStudy (zarkon):

yep...good stuff!

myininaya (myininaya):

i do love using the definition of derivative though

OpenStudy (anonymous):

next semester... if you start with \[fg=\frac{1}{2}((f+g)^2-f^2-g^2)\] puts all the computation on one side of the equal sign, easier for students to follow usually

myininaya (myininaya):

no it won't help lol nothing will

OpenStudy (anonymous):

ask them do it themselves. the algebra i mean.

OpenStudy (zarkon):

lol

myininaya (myininaya):

i already give up them and haven't even met them

myininaya (myininaya):

give up on them*

myininaya (myininaya):

i might ask my cal 2 students to do it on a test

myininaya (myininaya):

they don't even know how to differentiate so it will be good for them

OpenStudy (anonymous):

now you sound like a math teacher...

OpenStudy (anonymous):

it is generally good to assume little or no knowledge and start from scratch. at least where i am

myininaya (myininaya):

i thought i sounded like an english teacher

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

try 3 one hour twenty minute classes in a row! not enough americanos in the whole coffee shop to get me through